A professor needs to assign 15 teaching assistants to the course STATISTICS 101. There are 40 teaching assistants available for selection. (a) Since the professor doesn’t know any of the teaching assistants, they decide to take a random sample of 15 teaching assistants. How many different samples are possible? (b) Of the 40 available teaching assistants, 24 of them have experience with STATISTICS 101. What is the probability that all 15 of the teaching assistants selected by the professor have experience with this course?
Answer #1

Total assistants = 40

Total assistants to select = 15

Since order of selection is not important, we will use combination.

a) Number of different samples = 40C15 = 40! / (15! * (40 - 15)!) = 40225345056

b) Number of ways to select 15 teaching assistants with experience = 24C15 = 24! / (15! * (24 - 15)!) = 1307504

P(all 15 of the teaching assistants selected by the professor have experience with this course) = 1307504 / 40225345056 = 253 / 7783542